# Lower tensor-to-scalar ratio as possible signature of modified gravity

### v1

@@ I Introduction
As largely accepted, inflation became one of the cornerstone of modern cosmology. It not only solves the flatness and horizon problems, but also states the quantum seeds of cosmological fluctuations that eventually drove the universe to evolve [1, 2, 3, 4, 5, 6, 7, 8, 9] throughout a brief and rapid exponential expansion of the universe right after Big Bang. This inflationary period is thought to have smoothed out the early universe’s irregularities and laid the foundation for the large-scale structure we observe today. It is reinforced by the anisotropy seen in Cosmic Microwave Background (CMB) observations [10]. While the precise mechanisms fueling inflation remain elusive, numerous competing theoretical models have emerged [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] , each vying to explain this fundamental cosmic process. However, despite the ongoing debate, the overarching framework consistently yields predictions that align remarkably well with cosmological observations (see [22, 23] for a review).
@@
In this paper, we propose a possibility of the extrinsic curvature works as an inflaton field. Differently from traditional extra-dimensional models [32, 33, 34, 35], the extrinsic curvature is thought as the orthogonal component of the gravitational field besides the metric $g_{\mu\nu}$ as its tangent component. Our leitmotiv focus on how a dynamic embedding should affect the physical universe and may provide new insights on the current problems in cosmology. As a main character in this framework, the extrinsic curvature should not be restricted to the analysis of the background geometry but it should also include its perturbative dynamics [36, 37, 38]. To do so, we adopt Nash-Greene theorem [39, 40] that provides a general structure for embedding between non-Riemannian geometries [41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. It worth noting that, we do not use specific conditions on the extrinsic curvature such as Israel-Darmois-Lanczos (IDL) condition [51]. Commonly used in Randall-Sundrum models [33, 34], IDL condition replaces the extrinsic curvature by an algebraic relation with the energy-momentum tensor. It was showed it belongs to a very special case [52, 42], and, in general, IDL condition can be completely removed or replaced [53].
The paper is organized in sections. The second section verses on the essentials of embeddings as a mathematical background and summarises our cosmological model with both background and perturbed cosmological equations as well as the effective fluid approach necessary to develop a more realistic model to compare with observations. In the third section, we present a simpler version of the model in the form of a Modified Gravity (MG) framework. The fourth section is devoted to the analysis of extrinsic curvature as an inflaton field in which we investigate the slow-roll conditions for the extrinsic inflation potential. In sections V and IV, we examine the model in contrast with cosmological data with the latest NPIPE Planck DR4 likelihoods [54, 55], the BICEP2/Keck collaboration [56] and a junction of Large-scale structure (LSS) catalogue with 6dF Galaxy Survey [57], the Seventh Data Release of SDSS Main Galaxy Sample (SDSS DR7 MGS) [58] and clustering measurements of the Extended Baryon Oscillation Spectroscopic Survey (eBOSS) associated with the SDSS’s Sixteenth Data Release [59]. The related joint likelihood analysis is computed with MGCAMB-II [60] patch by means of Cobaya [61] sampler. In the final section, we conclude with our remarks and prospects.
@@ II Embeddings as a theoretical background
$k_{\mu\nu}=-\mathcal{X}^{A}_{,\mu}\;\eta^{B}_{,\nu}\mathcal{G}_{AB}\;.$ (2)
Eq.(2) shows the projection of the variation of normal unitary vector $\eta^{B}$ onto the tangent plane. The embedding coordinate $\mathcal{X}$ defines a regular local map $\mathcal{X}:V_{4}\rightarrow V_{5}$. The bulk metric $\mathcal{G}_{AB}$ is defined as
@@
Concerning notation, capital Latin indices run from 1 to 5. Overbar terms are background quantities unless. Small case Latin indices refer to the only one extra dimension considered. All Greek indices refer to the embedded space-time counting from 1 to 4.
@@
where $y$ is an arbitrary spatial coordinate. As a consequence of Eq.(6), we have how new geometries $g_{\mu\nu}$ are generated from small perturbations on the background metric increments $\delta g_{\mu\nu}$ as
@@
where $\delta k_{\mu\nu}=-2\delta y\;g^{\sigma\rho}\;k_{\mu\sigma}k_{\nu\rho}$. In principle, this is a continuous process to an any arbitrary perturbation increments of superior orders of $\delta g_{\mu\nu}$ and $\delta k_{\mu\nu}$. For our purposes, we keep the perturbations at linear order as show in Eq.(7) and (8).
The increments $\delta g_{\mu\nu}$ and $\delta k_{\mu\nu}$ are the Nash’s fluctuations later applied to non-Riemannian metrics by Greene [40]. The deformation formula Eq.(6) is also a pivot element to obtain solutions of the Gauss and Codazzi equations
@@
where ${}^{5}{\cal R}_{ABCD}$ is the five-dimensional Riemann tensor and $\bar{R}_{\alpha\beta\gamma\delta}$ is the background four-dimensional Riemann tensor. The perturbed embedding coordinate $\mathcal{Z}^{A}_{,\mu}$ is defined as $\mathcal{Z}^{A}_{,\mu}=\mathcal{X}^{A}_{,\mu}+\delta y\;\eta^{A}_{,\mu}\;$. The normal vector $\eta^{A}$ is invariant under perturbations, i.e., $\eta^{A}=\bar{\eta}^{A}$. The semicolon in Eq.(10) denotes covariant derivative with respect to the metric and the brackets apply the covariant derivatives to the adjoining indices.
@@ II.1 Background cosmology
where the scale factor is denoted by $a\equiv a(t)$ and $t$ is the physical time. Once the embedding is properly set, we impose that the bulk geometry is a solution of Einstein’s equations. Thus, from Eq.(9) and  Eq.(10), one obtains the non-perturbed field equations as
@@
where $\bar{T}_{\mu\nu}$ is the non-perturbed energy-momentum tensor of the confined perfect fluid and $G$ is the gravitational Newtonian constant. The background tensors $\bar{G}_{\mu\nu}$ and $\bar{Q}_{\mu\nu}$ represent the four dimensional Einstein tensor and the extrinsic deformation tensor, respectively. For an arbitrary D-dimensional case, see the detail derivation of Eq.(12) and Eq.(13[42] .
We can define the non-perturbed $\bar{Q}_{\mu\nu}$ in Eq.(12) is given by
@@
where $h^{2}=h\!\cdot\!h$ denotes the mean curvature, being $h=\;\bar{g}^{\mu\nu}\;\bar{k}_{\mu\nu}$. The Gaussian curvature is denoted by $K^{2}=\bar{k}^{\mu\nu}\bar{k}_{\mu\nu}$. A direct consequence of the previous definition in Eq.(14) is that the deformation tensor $\bar{Q}_{\mu\nu}$ is a conserved quantity such as
@@
As an FLRW solution, the trace of Codazzi equation of Eq.(13) and Eq.(14) were found as [42]
@@
where $\bar{Q}$ denotes the deformation scalar defined in a standard way, i.e., by the contraction $\bar{g}^{\mu\nu}\bar{Q}_{\mu\nu}=\bar{Q}$. One important definition is the evolution of the bending function $b(t)=b\equiv b(t)=k_{11}=k_{22}=k_{33}$ driven by extrinsic geometry. Thus, we defined $B=B(t)\equiv\frac{\dot{b}}{b}=(db/dt)/b$ as a copy of the Hubble parameter $H\equiv H(t)=\frac{\dot{a}}{a}=(da/dt)/a$. The $B(t)$ function has the same unit as $H$.
As shown in [42, 44], one obtains
@@ II.2 Perturbations in Conformal Newtonian gauge
The perturbed equations are necessary for the right estimation of cosmological parameters. In this framework, the relevant modifications of field equations under cosmological perturbations are applied to LHS of Eq.(12) with the presence of $\bar{Q}_{\mu\nu}$, and $\bar{G}_{\mu\nu}$ and $\bar{T}_{\mu\nu}$ tensors are treated in a very standard fashion as shown in Ref. [44]. Thus, a fluid with pressure $P$ and density $\bar{\rho}$, the perturbed components of the stress-tensor energy $\delta T_{\mu\nu}$ are given as
@@
where the prime symbol ${}^{\prime}$ denotes the ordinary derivative with respect to scale factor as ${}^{\prime}=\frac{d}{da}$.
@@
Just for notation sake and using Eq.(25), in the foregoing set of equations $\gamma_{0}$ denotes
@@
where $\beta_{0}$ was simplified by the introduction of the $\gamma_{s}$ parameter that is regarded as a relic from extrinsic geometry to maintain the characteristic of the parameter $\gamma_{0}$. Moreover, it turns $\gamma_{0}$ in Eq.(54) independent of the nature of the fluid characterized by $w$ parameter, as expected. This is also necessary to maintain the RG correspondence that is achieved when extrinsic curvature vanishes, i.e., $\gamma_{s}\rightarrow 0$.

### v2

# Lower tensor-to-scalar ratio as a possible signature of modified gravity

@@ I Introduction
As largely accepted, inflation became one of the cornerstone of modern cosmology. It does not only solve the flatness and horizon problems, but also states the quantum seeds of cosmological fluctuations that eventually drove the universe to evolve [1, 2, 3, 4, 5, 6, 7, 8, 9] throughout a brief and rapid exponential expansion of the universe right after Big Bang. This inflationary period is thought to have smoothed out the early universe’s irregularities and laid the foundation for the large-scale structure we observe today. It is reinforced by the anisotropy seen in Cosmic Microwave Background (CMB) observations [10]. While the precise mechanisms fueling inflation remain elusive, numerous competing theoretical models have emerged [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], each vying to explain this fundamental cosmic process. However, despite the ongoing debate, the overarching framework consistently yields predictions that align remarkably well with cosmological observations (see [22, 23] for a review).
@@
In this paper, we propose the possibility for the extrinsic curvature, thought as the orthogonal component of the gravitational field besides the metric $g_{\mu\nu}$, works as an inflaton field. Differently from traditional extra-dimensional braneworld models [32, 33, 34, 35], the dynamics of extrinsic curvature is considered. Our leitmotiv focus on how a dynamic embedding should affect the physical universe and may provide new insights on the current problems in cosmology. As a main character in this framework, the extrinsic curvature should not be restricted to the analysis of the background geometry but it should also include its perturbative dynamics [36, 37, 38]. To do so, we start from the perturbations of the geometry adopting Nash-Greene theorem [39, 40] that provides a general structure for embedding between non-Riemannian geometries [41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. This is rather different from common approaches of brane-world perturbative models in which the perturbations are triggered from confined sources. Then the dynamics of the extrinsic curvature itself is replaced. A common practice in these models is to rely on junction conditions such as Israel-Darmois-Lanczos (IDL) condition [51]. Commonly used in Randall-Sundrum models [33, 34], IDL condition replaces the extrinsic curvature by an algebraic relation with the energy-momentum tensor. It was showed it belongs to a very special case [52, 42], and, in general, IDL condition can be completely removed or replaced [53].
The paper is organized in sections. The second section verses on the essentials of embeddings as a mathematical background and summarizes our cosmological model with both background and perturbed cosmological equations. Moreover, the effective fluid approach is presented to develop a more realistic model to compare with observations. In the third section, we present a simpler version of the model in the form of a Modified Gravity (MG) framework. The fourth section is devoted to the analysis of extrinsic curvature as an inflaton field in which we investigate the slow-roll conditions for the extrinsic potential. In sections V and IV, we examine the model in contrast with cosmological data with the latest NPIPE Planck DR4 likelihoods [54, 55], the BICEP2/Keck collaboration [56] and a junction of Large-scale structure (LSS) catalogue with 6dF Galaxy Survey [57], the Seventh Data Release of SDSS Main Galaxy Sample (SDSS DR7 MGS) [58] and clustering measurements of the Extended Baryon Oscillation Spectroscopic Survey (eBOSS) associated with the SDSS’s Sixteenth Data Release [59]. The related joint likelihood analysis is computed with MGCAMB-II [60] patch by means of Cobaya [61] sampler. In the final section, we conclude with our remarks and prospects.
@@ II Embeddings as a theoretical background
$\bar{k}_{\mu\nu}=-\mathcal{X}^{A}_{,\mu}\;\bar{\eta}^{B}_{,\nu}\mathcal{G}_{AB}\;,$ (2)
where Eq.(2) shows the projection of the variation of the set of normal unitary vector $\bar{\eta}^{B}$ onto the tangent plane orthogonal to the embedded space $V_{4}$. In other words, the variation of $\bar{\eta}^{B}$ leads to the bending of $V_{4}$ and it has its tangent components with coefficients $\bar{k}_{\mu\nu}$. The embedding coordinate $\mathcal{X}$ defines a regular local map $\mathcal{X}:V_{4}\rightarrow V_{5}$ and must satisfy the embedding equations
$\mathcal{X}^{A}_{,\mu}\mathcal{X}^{B}_{,\nu}\mathcal{G}_{AB}=g_{\mu\nu},\;\mathcal{X}^{A}_{,\mu}\bar{\eta}^{B}\mathcal{G}_{AB}=0,\;\bar{\eta}^{A}\bar{\eta}^{B}\mathcal{G}_{AB}=1,$ (3)
where we have denoted by $\mathcal{G}_{AB}$ the metric components of the bulk $V_{5}$ in arbitrary coordinates. The embedding frame is defined by the set of coordinates $\{\mathcal{X}^{A},\bar{\eta}^{A}\}$ that composes a Gaussian reference frame. Throughout the paper, except when explicitly stated in contrary, the overbar sign indicates a background (non-perturbed) quantity.
The bulk metric $\mathcal{G}_{AB}$ is defined as
@@
Concerning notation, capital Latin indices run from 1 to 5. Small case Latin indices refer to the only one extra dimension considered. All Greek indices refer to the embedded space-time counting from 1 to 4.
@@
where $y$ is an arbitrary spatial coordinate along the orthogonal direction to the tangent plane. This mechanism avoids false perturbations due to the possibility to induce coordinate gauges. Therefore, the physical effects of the extrinsic curvature associated with Eq.(7) represent an acceleration tangent to the four-dimensional space-time that always points to the concave side of the curve. As a result, it induces a Riemann stretching on the space-time geometry, which may be related to the accelerated expansion of the universe [42, 44].
As a consequence of Eq.(7), we have how new geometries $g_{\mu\nu}$ are generated from small perturbations on the background metric increments $\delta g_{\mu\nu}$ as
@@
where $\delta k_{\mu\nu}=-2\delta y~{}g^{\sigma\rho}k_{\mu\sigma}k_{\nu\rho}$. In principle, this is a continuous process to any arbitrary perturbation increments of superior orders of $\delta g_{\mu\nu}$ and $\delta k_{\mu\nu}$. For our purposes, we keep the perturbations at linear order as show in Eq.(8) and (9).
The increments $\delta g_{\mu\nu}$ and $\delta k_{\mu\nu}$ are the Nash’s fluctuations later applied to non-Riemannian metrics by Greene [40]. The deformation formula in Eq.(7) is also a pivot element to obtain solutions of the Gauss and Codazzi equations
@@
where ${}^{5}{\cal R}_{ABCD}$ is the five-dimensional Riemann tensor and $\bar{R}_{\alpha\beta\gamma\delta}$ is the background four-dimensional Riemann tensor. The perturbed embedding coordinate $\mathcal{Z}^{A}_{,\mu}$ is defined as $\mathcal{Z}^{A}_{,\mu}=\mathcal{X}^{A}_{,\mu}+\delta y\;\eta^{A}_{,\mu}\;$. The normal vector $\eta^{A}$ is invariant under perturbations, i.e., $\eta^{A}=\bar{\eta}^{A}$. The semicolon in Eq.(11) denotes covariant derivative with respect to the metric and the brackets apply the covariant derivatives to the adjoining indices. In Eq.(10), the Gauss equation entwines the bulk Riemann curvature as a reference standpoint for the Riemann curvature of the embedded space-time. This relation is complemented by the Codazzi equation in Eq.(11), which shows the projection of the Riemann tensor of the bulk space along the orthogonal direction results in the variation of the extrinsic curvature.
In order to guarantee the possibility to generate new geometries, in five-dimensions, the set of coordinates $\mathcal{Z}^{A}$ also needs to satisfy embedding equations similar to Eq.(3):
$\mathcal{Z}^{A}{}_{,\mu}\mathcal{Z}^{B}{}_{,\nu}\mathcal{G}_{AB}=g_{\mu\nu},\;\mathcal{Z}^{A}{}_{,\mu}\eta^{B}\mathcal{G}_{AB}=0,\;\eta^{A}\eta^{B}\mathcal{G}_{AB}=1\;,$ (12)
then, Eq.(8) and (9) are valid.
In a similar fashion as Kaluza-Klein and brane-world models, we consider that the bulk $V_{5}$ has its dynamics governed by the higher-dimensional Einstein’s equations
${}^{5}{\mathcal{R}}_{AB}-\frac{1}{2}\,{{}^{5}{\mathcal{R}}}{\mathcal{G}}_{AB}=G_{*}T^{*}_{AB}$ (13)
where $G_{*}$ is the new gravitational constant and $T^{*}_{AB}$ denotes the components of the energy-momentum tensor of the material sources. Those sources are confined to four dimensions. It is a consequence of the isomorphism between the three-form, from the derivative of the Yang-Mills curvature, and one-form current. This is valid only in four dimensions. Thus, all known observable sources of gravitation composing the generic energy-momentum tensor $T^{*}_{AB}$ are also confined. This outcome is independent of the variation of extra coordinate $y$. Otherwise speaking, the four-dimensionality of space-time is a consequence of the invariance of Maxwell’s equations under the Poincaré group. It is well known that any gauge theory can be mathematically constructed or extended in a higher dimensional space, but in the present framework, the four-dimensionality of the embedded space-time will suffice based on what experimentally high-energy tests suggest [63, 64, 65]. Moreover, in the Gaussian frame $\{\mathcal{Z}^{A},\eta^{A}\}$ of any perturbed space-time, we can write the confinement condition of the energy-momentum tensor source $T^{*}_{AB}$ of the bulk Einstein’s equation in Eq.(13) with the projections
$8\pi GT_{\mu\nu}=G_{*}\mathcal{Z}^{A}_{,\mu}\mathcal{Z}^{B}_{,\nu}T^{*}_{AB},\mathcal{Z}^{A}_{,\mu}\eta^{B}T^{*}_{AB}=0,\eta^{A}\eta^{B}T^{*}_{AB}=0,$ (14)
In this framework, the matter content is localized in the $V_{4}$ embedded space due to the fact that Nash’s deformation formula in Eq.(7) imposes a geometric constraint on the confined sources. Any deformation is not arbitrary and can be generated by smooth perturbations along the direction $\delta y$ orthogonal to the embedded space $V_{4}$. In five dimensions, this process is simplified and just one deformation parameter suffices to locally deform the embedded background. In general, the curvature radii $l^{a}$ of the embedded background correspond to the direction in which the embedded space-time deviates more sharply from the tangent plane and are the solutions of the homogeneous equation
$\mbox{det}(g_{\mu\nu}-l^{a}k_{\mu\nu a})=0\;.$ (15)
This is a local invariant property of the embedded space-time and does not depend on the chosen Gaussian system. The smallest solution $l$ provides
$\frac{1}{l}=\left(g_{\mu\nu}\mathcal{G}^{AB}l_{A}^{\mu}l_{B}^{\nu}\right)^{-1/2}\;.$ (16)
Otherwise speaking, $l$ constrains a local limit for the region in which the bulk is accessed by the gravitons. Then, one can find the typical length $d$ [41] of the n-extra dimensional space accessed by gravitons as
$d=\frac{M_{Pl}^{2/n}}{M_{\ast}^{1+(2/n)}}\frac{1}{(1+\frac{M_{Pl}^{2}}{M_{e}^{2}})^{1/n}}\;,$ (17)
where $M_{\ast}$ and $M_{Pl}$ are respectively the fundamental and effective Planck scales. The extrinsic scale $M_{e}$ is given by
$\frac{1}{M_{e}}=\int(K^{2}+h^{2})\sqrt{g}d^{4}x\;,$ (18)
where $K^{2}$ is the Gaussian curvature and $h^{2}$ is the mean curvature. Hence, for smooth oscillations of the embedded background, the limit imposed by Eq.(17) with the $M_{e}$ scale, prevents the leak of energy of the confined sources to higher dimensional space but it allows the graviton oscillations. This eliminates the necessity of introduction of a radion field, e.g., commonly adopted in most popular braneworld models [32, 33, 34, 35].
@@ II.1 Background cosmology
where the scale factor is denoted by $a\equiv a(t)$ and $t$ is the physical time. Once the embedding is properly set, we impose that the bulk geometry is a solution of Einstein’s equations given by Eq.(13). Thus, from Eq.(10) and  Eq.(11), and the confinement condition in Eq.(14), one obtains the tangent components of non-perturbed field equations as
@@
where $\bar{T}_{\mu\nu}$ is the non-perturbed energy-momentum tensor of the confined perfect fluid and $G$ is the gravitational Newtonian constant. The background tensors $\bar{G}_{\mu\nu}$ and $\bar{Q}_{\mu\nu}$ represent the four dimensional Einstein tensor and the extrinsic deformation tensor, respectively. For an arbitrary D-dimensional case, see the detail derivation of Eq.(20) and Eq.(21) in Ref. [42].
In addition, the non-perturbed $\bar{Q}_{\mu\nu}$ in Eq.(20) is defined as
@@
where $h^{2}=h\!\cdot\!h$ denotes the mean curvature with $h=\;\bar{g}^{\mu\nu}\;\bar{k}_{\mu\nu}$. The Gaussian curvature is denoted by $K^{2}=\bar{k}^{\mu\nu}\bar{k}_{\mu\nu}$. A direct consequence of the previous definition in Eq.(22) is that the deformation tensor $\bar{Q}_{\mu\nu}$ is a conserved quantity such as
@@
Solving the trace of Codazzi equation in Eq.(21) and Eq.(22), the following components were found [42]
@@
where $\bar{Q}$ denotes the deformation scalar defined in a standard way, i.e., by the contraction $\bar{g}^{\mu\nu}\bar{Q}_{\mu\nu}=\bar{Q}$. One important definition is the evolution of the bending function $b(t)\equiv b=k_{11}=k_{22}=k_{33}$ driven by extrinsic geometry. Thus, we defined $B=B(t)\equiv\frac{\dot{b}}{b}=(db/dt)/b$ as a copy of the Hubble parameter $H\equiv H(t)=\frac{\dot{a}}{a}=(da/dt)/a$. As a consequence, the $B(t)$ function borrows the same unit of $H$.
An important aspect is that in five dimensions, the trace of Codazzi equation in Eq.(21) is homogeneous which turns the solution for the bending function $b(t)$ arbitrary. To remove such arbitrariness, we need to state the dynamics of the extrinsic curvature, since the metric and the extrinsic curvature are independent variables that must satisfy the Gauss (Eq.(10)) and Codazzi (Eq.(11)) equations. Thus, one counts a total of 20 unknowns $g_{\mu\nu}$ and $k_{\mu\nu}$, against only 15 dynamical equations. This requires to consider $k_{\mu\nu}$ as a source for the missing equations. A well known theorem due to S. Gupta [66] states that any symmetric rank-2 tensor satisfies an Einstein-like system of equations, having the Pauli-Fierz equation as its linear approximation. As proposed by Isham, Salam, and Strathdee [67] in the context of strong gravity, they propose a f-meson spin-2 field that would act as an intermediate field between gravitation and hadron particles. In a previous publication [44], an Einstein-like dynamical equation for the extrinsic curvature was adapted from the original equation of S. Gupta. Using Gupta’s theorem, as shown in [44, 49], one simply obtains
@@ II.2 Perturbations in Conformal Newtonian gauge
The perturbed equations are necessary for the right estimation of cosmological parameters. In this framework, the relevant modifications of field equations under cosmological perturbations are applied to LHS of Eq.(20) with the presence of $\bar{Q}_{\mu\nu}$, and $\bar{G}_{\mu\nu}$ and $\bar{T}_{\mu\nu}$ tensors that are treated in a very standard fashion as shown in Ref. [44]. Thus, a fluid with pressure $P$ and density $\bar{\rho}$, the perturbed components of the stress-tensor energy $\delta T_{\mu\nu}$ are given as
@@
where the prime symbol denotes the ordinary derivative with respect to scale factor as ${}^{\prime}=\frac{d}{da}$.
@@
For the sake of notation and using Eq.(33) in the foregoing set of equations, $\gamma_{0}$ denotes
@@
where $\beta_{0}$ was simplified by the introduction of the $\gamma_{s}$ parameter that is regarded as a relic from extrinsic geometry to maintain the characteristic of the parameter $\gamma_{0}$. Moreover, it turns $\gamma_{0}$ in Eq.(62) independent of the nature of the fluid characterized by $w$ parameter, as expected. This is also necessary to maintain the RG correspondence that is reached when extrinsic curvature vanishes, i.e., $\gamma_{s}\rightarrow 0$.